OpenStudy (fluffynose):

MEDALS MEDALS MEDASLS fill in the blanks please. A circle is centered at the point (-3, 2) and passes through the point (1, 5). The radius of the circle is________ units. The point (-7, _____ ) lies on this circle.

OpenStudy (xx_gamergirl_xx):

Are there any choices for the words?

OpenStudy (fluffynose):

nooo. :(

OpenStudy (xx_gamergirl_xx):

Okay, so you know a radius means the length between the middle and one side, and the diameter is the length from one side to the other.

OpenStudy (fluffynose):

Yes but i dunno how to figure out?

OpenStudy (xx_gamergirl_xx):

So knowing that, and knowing the center of the circle, (-3,2) and knowing where is ends, (1,5) how long would that be from (-3,2) to (1,5)?

rishavraj (rishavraj):

http://prntscr.com/agu3qo apply distacne formula and get the radius

OpenStudy (fluffynose):

2,3?

OpenStudy (fluffynose):

@xX_GamerGirl_Xx

OpenStudy (fluffynose):

@Keigh2015

OpenStudy (keigh2015):

The radius I believe would just be one number.

OpenStudy (faiqraees):

Use the formula \[\large\rm (y-y_1)^2+(x-x_1)^2=r^2 \] where (x1,y1) are the centre of the circle

OpenStudy (fluffynose):

okay i got 25 and 5? are those rght?

rishavraj (rishavraj):

@fluffynose the pic i posted ...accordng to it (x1, y1) = (-3 , 2) and (x2 , y2) = (1, 5) now apply the distance formual to get OP ...i.e the radius \[OP = \sqrt{(-3-1)^2 + (2 - 5)^2}\]

rishavraj (rishavraj):

remember OP is the radius :)

rishavraj (rishavraj):

once u get the radius u can solve the 2nd part easily :)) @fluffynose

rishavraj (rishavraj):

for 2nd part :)) OP = OQ = radius http://prntscr.com/agug3j

rishavraj (rishavraj):

@Keigh2015 yup thts correct :)